# Mix Q-learning for Lane Changing: A Collaborative Decision-Making Method in Multi-Agent Deep Reinforcement Learning

### v1

@@ III-B Deep Reinforcement Learning Algorithm
Figure 2: The overall framework of MQLC. The black solid lines represents the original decision process, the green solid lines represents the additional improvements to the decision process in MQLC, and the blue and red dashed lines represent the process of collecting the surrounding vehicle trajectories. The decision making process of the intelligent agent can be described as follows: first, MQLC estimates the intentions of the surrounding vehicles based on their trajectories. This information is then fused with the original observations and collectively used for decision making. Next, at the individual level, the intelligent agent uses an innovative network structure to optimise information embedding and estimate feasible decisions. Finally, at the global level, an advantage estimation is performed for the joint actions that all intelligent agents can take. This estimation guides the intelligent agents in achieving collaborative lane-changing decisions.
@@
In order to achieve true collaboration in lane-changing decisions, we referenced Zhang’s QCOMBO[30] algorithm proposed for signal light optimisation when designing the algorithm framework. We then made modifications to adapt it to the characteristics of the lane-changing task. QCOMBO is a value-based reinforcement learning network. Compared to DQN, it combines centralised learning and independent learning, which is manifested in its individual Q and global Q. The optimisation objective of Qcombo consists of three parts. In addition to the losses from the two reward-feedback networks, an additional regulariser is introduced to minimise the difference between the weighted sum of the global Q and the individual Q. The specific design of the losses is as follows:
(a) Individual part: The training goal for the individual part is to gradually make the individual Q network’s advantage estimates for each action in the observation approach the true feedback reward, based on the agent’s interactions in the environment. To improve the training efficiency and stability, a strategy of using local reward[27] and parameter sharing[39, 40] is adopted in this part. In terms of lane-changing decisions, agents share the same observation and action space. By using parameter sharing in an execution, the environment can obtain multiple data sets. This allows for a richer data set to train the Q-network, thereby improving training efficiency.
For the individual Q-network of agent $i$ parameterised by $\theta^{i}$, we minimise the following loss
$\mathcal{L}_{ind}(\theta^{i})=\mathbb{E}_{\left(o_{t}^{i},a_{t}^{i},r_{t}^{i},o_{t+1}^{i}\right)\sim D_{ind}}\left[\left(y_{t}^{i}-Q_{i}\left(o_{t}^{i},a_{t}^{i};\theta^{i}\right)\right)^{2}\right],\$ (6)
@@
(b) Global part: The purpose of the global part is to train a global Q network that can provide estimates of the benefits of joint actions based on global information and come close to the total reward feedback. Due to the necessity of selecting the best action out of n possible joint actions during the learning process of Q networks, the design of the global Q is limited in scalability in many multi-agent tasks. To address this issue, QCOMBO proposes the idea of training the global Q using local utility functions. The global Q treats the action taken by each agent’s individual Q when it is optimal as the joint action and only estimates the utility of that action. Specifically, the joint action is chosen by $\mathbf{a}\sim\boldsymbol{\pi}(\mathbf{a}\mid s)=\left\{\operatorname{argmax}_{a^{i}}Q^{n}\left(o^{i},a^{i}\right)\right\}_{i=1}^{N}$, and its value is defined by $R_{t}^{g}:=R_{t}^{g}(s,\mathbf{a})$. Based on the above, the global Q can be written as
@@
For the global Q-network parameterised with $\omega$, we minimise the following loss
@@
To address these limitations, we introduce an innovative decision-making process in the MQLC framework, where coordination is achieved through the collaborative efforts of two-level Q-networks. Specifically, in a single decision instance within the MQLC framework, each agent first uses its individual Q to evaluate the benefits of different actions based on its own observations. To ensure the generalisability and reliability of MQLC, we introduce an additional urgency concept, i.e. the decision priority, by considering the surrounding traffic conditions. The decision priority is defined by the following formula
@@
where $priority_{i}$ is the decision priority of the agent $i$, $urgency_{i}$ is the quantified level of congestion, $m\_velocity$ is the average speed of the agent and the surrounding vehicles, $t\_density$ is the traffic density in the environment surrounding the agent, while s_variance is the variance in speeds between the agent and the surrounding vehicles. $\alpha$ is the weight of the speed variance term. In this way, MQLC can adapt to different complex traffic scenarios by simply adjusting the urgency parameters, without having to repeat complex reinforcement learning training. At the same time, the priority concept, which is independent of the Q-network, can ensure the basic interests of the agent, rather than completely obeying the whole for the highest global reward.
@@
2:     $//Collect\ training\ data$
3:     for $Vehicle\ i=1$ to $N$ do
@@ III-B2 The specific design of the value network
The two value networks we propose are both based on deep neural networks. The individual Q network takes the observation $o$ as input and outputs payoff estimates for individual actions. On the other hand, the global Q network takes the state $s$ as network input and outputs benefit estimates for joint actions that can be taken. Given the matrix input form, both networks use Multi-Layer Perceptron (MLP) as their main components. MLP uses its strong feature extraction and generalisation capabilities to fulfil the specific functions of value networks.
Although MLP is widely used in decision learning, it has inadequate information extraction capability limitations, which leads to less efficient decision making.
To overcome these limitations, MQLC uses a combination of several network architectures in the decision making process. The MQLC decision network architecture is shown in Figure 3.
@@ IV Performance evaluation
In this section, we perform experiments and analyses of the proposed method. Subsection IV-A introduces the simulation environment and the parameter settings in the experiments. Subsection IV-B presents the result of the comparative experiments. Subsection IV-C discusses the parameter settings of the training state. Section IV-D performs ablation experiments to explore the role of different components of the model. Subsection IV-E conducts qualitative experiments to investigate the effect of cooperative strategy and emergency situation settings on decision making.
@@ IV-A Simulation environment and parameters settings
We constructed a 1000m six-lane motorway environment for training and testing. All vehicles were initialised to random lanes with a fixed interval and were travelling at speeds between 20 $km/h$ and 30 $km/h$ at the beginning of the test. The proposed lane-changing decision model was built in the above environment. Due to the similarity in the format of observations and states in this task, both individual Q and global Q adopted the same network architecture, as shown in Table II. The parameter settings of the environment and lane change model are shown in Table III.
@@
GCN Layer 1 Input size = $V$×2, Output size = 256, activation = $ReLU$
GCN Layer 2 Input size = 256, Output size = 256, activation = $ReLU$
FC Layer 1 Input size = 256, Output size = 256, activation = $ReLU$
GRU Layer Input size = $T_{P}$×256, Output size = 256, activation = $ReLU$
FC Layer 2 Input size = 256, Output size = 10, activation = $ReLU$
@@
GCN Layer 1 for $h_{sur}$ Input size = $V$×2(individual Q) or $N$×$V$×2(global Q), Output size = 256, activation = $ReLU$
GCN Layer 2 for $h_{sur}$ Input size = 256, Output size = 256, activation = $ReLU$
FC Layer 1 for $h_{sur}$ Input size = 256, Output size = 256, activation = $ReLU$
FC Layer 2 for $h_{env}$ Input size = 2, Output size = 8, activation = $ReLU$
FC Layer 3 for $h_{ori}$ Input size = $V$×$F$(individual Q) or $N$×$V$×$F$(global Q), Output size = 256, activation = $ReLU$
FC Layer 4 for $h_{ori}$ Input size = 256, Output size = 256, activation = $ReLU$
FC Layer 5 Input size = 520, Output size = 5(individual Q) or 125(global Q), activation = $ReLU$
@@
It is crucial to assess the impact of changing the function of global Q from ”estimating the value of the optimal joint action based on the state” to ”estimating the value of all joint actions based on the state” when configuring the parameters. If the global Q does not converge, it not only affects its judgement on joint actions, but also affects the training of the individual Q through the loss shown in Eq.(11), which may lead to a failure of decision making even at the individual level. During the training of the global Q, its convergence speed may be slower than that of the individual Q due to its decision-making on joint actions with larger action spaces and the limited data collected in each training iteration compared to the parameter-shared individual parts. To improve the overall learning efficiency, we adopt the approach of adjusting the learning rates and the dynamic learning rate [45] to balance the training of individual Q networks for decision making with the acquisition of global knowledge, while accelerating the convergence of the global Q network.
@@
IV-B Result of comparative experiments
To demonstrate the superiority of the proposed MQLC in decision making, we conduct comparative experiments. We emphasise the consideration of the reward component in the evaluation metrics, as rewards serve as the only feedback signal during reinforcement learning training. They directly reflect the quality of the actions taken by the intelligent agent.
@@ IV-B Result of comparative experiments
Based on the specific definition of the reward function proposed in chapter III-A4, the efficiency and safety of lane-changing decisions are also the key objectives we consider. The average speed of agents provides a clear understanding of the benefits of their decisions and can also indirectly reflect the overall traffic efficiency. We therefore choose it as a measure of efficiency. And in the simulation environment, collisions lead to the termination of episodes before the maximum episode length is reached. The longer the episode length is, the safer the lane-changing decision made by the agent is. As a result, the average episode length serves as an indicator for evaluating the safety of the models.
@@
Comparative experiments were carried out in the three traffic environments shown in Table IV. The ’sparse’ mode in the table represents a low-density traffic pattern where there are two agents and 8 surrounding vehicles; the ’normal’ mode represents a medium-density traffic pattern with 3 agents and 15 surrounding vehicles; and the ’dense’ mode represents a high-density traffic pattern with 5 agents and 30 surrounding vehicles. Table V shows the results of the comparative experiments in terms of safety, efficiency and reward metrics. From Table V, we can draw the following conclusions:
(1)Comparing the ’Random’ row in Table V with the other rows, it can be seen that all the reinforcement learning methods listed in the table are capable of acquiring some degree of lane-changing decision making ability after training, resulting in better lane-changing performance compared to ’Random’.
@@
(2)Comparing the ”MQLC” row in Table V with the ”DQN”, ”Double DQN” and ”D3QN” rows, it can be observed that MQLC achieves the best performance among the reinforcement learning methods. In the ”Normal” and ”Dense” traffic scenarios, MQLC achieves the highest average episode length, the highest speed, and the optimal reward. However, in the ’Sparse’ scenario, all models show similar performance and MQLC’s performance is not significantly better than the other models. This is because MQLC’s primary innovation lies in the coordination between multiple intelligent agents. When there are relatively few agents, MQLC does not show a clear advantage. However, as the number of agents increases, MQLC can optimise the overall traffic flow by using additional feature extraction and coordinated decision making, balancing individual and global interests, and thus achieving optimisation of the overall traffic flow.
@@
(3) Comparing the ”MQLC” row with the ”QCOMBO” row shows that MQLC consistently outperforms QCOMBO in various scenarios. This is because QCOMBO only considers bringing individual network estimates closer to the global network during training, but continues to use individual decisions during actual decision making, and is unable to overcome the limitations of information acquisition and decision coordination. In contrast, MQLC fully exploits the advantages of CVIS, uses state information to understand the overall traffic situation, and effectively coordinates during decision making by using individual agents to support global decision making. As a result, MQLC achieves significantly better lane-changing decision performance than QCOMBO.
@@
(4) Comparing the three performance metrics across various scenarios, MQLC consistently outperforms other models in terms of ”Length” and ”Reward” in nearly all cases. However, except for the most congested ”Dense” scenario, there isn’t a significant advantage in the ”Speed” metric in the other two scenarios. These results suggest that MQLC places a higher emphasis on prioritizing the safety of all agents’ actions through collaborative efforts, thereby avoiding collision penalties and ultimately achieving superior lane-changing decisions, rather than solely focusing on efficiency enhancement.
TABLE VI: The effect of $\lambda$ in performance

### v2

@@ III-B Deep Reinforcement Learning Algorithm
Figure 2: The overall framework of MQLC. The decision making process of the intelligent agent can be described as follows: first, MQLC estimates the intentions of the surrounding vehicles based on their trajectories. Next, at the individual level, the agent uses an innovative network structure to optimise information embedding and estimate feasible decisions. Finally, at the global level, an advantage estimation is performed for the joint actions that all intelligent agents can take.
@@
In order to achieve true collaboration in lane-changing decisions, we reference QCOMBO[30] algorithm proposed for signal light optimisation when designing the algorithm framework. We then made modifications to adapt it to the characteristics of the lane-changing task. QCOMBO is a value-based reinforcement learning network. Compared to DQN, it combines centralised learning and independent learning, which is manifested in its individual Q and global Q. The optimisation objective of QCOMBO consists of three parts. In addition to the losses from the two reward-feedback networks, an additional regulariser is introduced to minimise the difference between the weighted sum of the global Q and the individual Q. The specific design of the losses is as follows:
(a) Individual part: The training goal for the individual part is to gradually make the individual Q network’s advantage estimates for each action in the observation approach the true feedback reward, based on the agent’s interactions in the environment. To improve the training efficiency and stability, a strategy of using local reward[27] and parameter sharing[37, 38] is adopted in this part.
For the individual Q-network of agent $i$ parameterised by $\theta^{i}$, minimise the following loss
$\mathcal{L}_{ind}(\theta^{i})=\mathbb{E}_{\left(o_{t}^{i},a_{t}^{i},r_{t}^{i},o_{t+1}^{i}\right)\sim D_{ind}}\left[\left(y_{t}^{i}-Q_{i}\left(o_{t}^{i},a_{t}^{i};\theta^{i}\right)\right)^{2}\right],\\$ (3)
@@
(b) Global part: The purpose of the global part is to train a global Q network that can provide estimates of the benefits of joint actions based on global information and come close to the total reward feedback. The global Q treats the action taken by each agent’s individual Q when it is optimal as the joint action and only estimates the utility of that action. Specifically, the joint action is chosen by $\mathbf{a}\sim\boldsymbol{\pi}(\mathbf{a}\mid s)=\left\{\operatorname{argmax}_{a^{i}}Q^{n}\left(o^{i},a^{i}\right)\right\}_{i=1}^{N}$, and its value is defined by $R_{t}^{g}:=R_{t}^{g}(s,\mathbf{a})$. Based on the above, the global Q can be written as
@@
For the global Q-network parameterised with $\omega$, minimise the following loss
@@
To address these limitations, we introduce an innovative decision-making process in the MQLC framework, where coordination is achieved through the collaborative efforts of two-level Q-networks. To ensure the generalisability and reliability of MQLC, we introduce an additional urgency concept, i.e. the decision priority, by considering the surrounding traffic conditions. The decision priority is defined by the following formula,
@@
where $priority_{i}$ is the decision priority of the agent $i$, $urgency_{i}$ is the quantified level of congestion, m_velocitym\_velocity is the average speed of the agent and the surrounding vehicles, t_densityt\_density is the traffic density in the environment surrounding the agent, while s_variances\_variance is the variance in speeds between the agent and the surrounding vehicles. $\alpha$ is the weight of the speed variance term. In this way, MQLC can adapt to different complex traffic scenarios by simply adjusting the urgency parameters, without having to repeat complex reinforcement learning training. At the same time, the priority concept, which is independent of the Q-network, can ensure the basic interests of the agent, rather than completely obeying the whole for the highest global reward.
@@
In the MQLC method proposed in this article, there is no correlation between individual Q (-network) and global Q (-network) at the neural network architecture level, that is, there is no interpenetration or stacking between the layers of the two networks. The correlation between these two neural networks lies in the fact that the output of individual Q is a part of the input of global Q, which affects the output of global Q. Specifically, each agent determines the optimal and suboptimal actions that can be taken based on individual Q. The global Q selects one of the optimal and suboptimal actions that each agent can take based on the global situation, ultimately selecting the joint action that maximizes global expected benefits.
@@
2:  //Collecttrainingdata//Collect\ training\ data
3:  for Vehiclei=1Vehicle\ i=1 to NN do
@@ III-B2 The specific design of the value network
The two value networks we propose are both based on deep neural networks. The individual Q network takes the observation $o$ as input and outputs payoff estimates for individual actions. On the other hand, the global Q network takes the state $s$ as network input and outputs benefit estimates for joint actions that can be taken. Given the matrix input form, both networks use Multi-Layer Perceptron (MLP) as their main components. MLP uses its strong feature extraction and generalisation capabilities to fulfil the specific functions of value networks. Although MLP is widely used in decision learning, it has inadequate information extraction capability limitations, which leads to less efficient decision making. To overcome these limitations, MQLC uses a combination of several network architectures in the decision making process. The MQLC decision network architecture is shown in Figure 3.
@@ IV Performance evaluation
In this section, we perform experiments and analyses of the proposed method. Subsection IV-A introduces the simulation environment and the parameter settings in the experiments. Subsection IV-C presents the result of the comparative experiments. Subsection IV-E discusses the parameter settings of the training state. Section IV-F performs ablation experiments to explore the role of different components of the model. Subsection IV-G conducts qualitative experiments to investigate the effect of cooperative strategy and emergency situation settings on decision making.
@@ IV-A Simulation environment and parameters settings
We constructed a 1000m six-lane motorway environment for training and testing. All vehicles were initialised to random lanes with a fixed interval and were travelling at speeds between 20 $km/h$ and 30 $km/h$ at the beginning of the test. The proposed lane-changing decision model is built in the above environment. Due to the similarity in the format of observations and states in this task, both individual Q and global Q adopt the same network architecture, as shown in Table II. The parameter settings of the environment and lane change model are shown in Table III.
@@
GCN Layer 1 Input size = VV×2, Output size = 256, activation = ReLUReLU
GCN Layer 2 Input size = 256, Output size = 256, activation = ReLUReLU
FC Layer 1 Input size = 256, Output size = 256, activation = ReLUReLU
GRU Layer Input size = $T_{P}$×256, Output size = 256, activation = ReLUReLU
FC Layer 2 Input size = 256, Output size = 10, activation = ReLUReLU
@@
GCN Layer 1
for $h_{sur}$
Input size = VV×2(individual Q) or NN×VV×2(global Q),
Output size = 256, activation = ReLUReLU
GCN Layer 2
for $h_{sur}$
Input size = 256, Output size = 256,
activation = ReLUReLU
FC Layer 1
for $h_{sur}$
Input size = 256, Output size = 256,
activation = ReLUReLU
FC Layer 2
for $h_{env}$
Input size = 2, Output size = 8,
activation = ReLUReLU
FC Layer 3
for $h_{ori}$
Input size = VV×FF(individual Q) or NN×VV×FF(global Q),
Output size = 256, activation = ReLUReLU
FC Layer 4
for $h_{ori}$
Input size = 256, Output size = 256,
activation = ReLUReLU
FC Layer 5
Input size = 520, Output size = 5(individual Q)
or 125(global Q), activation = ReLUReLU
@@
It is crucial to assess the impact of changing the function of global Q from “estimating the value of the optimal joint action based on the state” to “estimating the value of all joint actions based on the state” when configuring the parameters. If the global Q does not converge, it not only affects its judgement on joint actions, but also affects the training of the individual Q through the loss shown in Eq.(8), which may lead to a failure of decision making even at the individual level. During the training of the global Q, its convergence speed may be slower than that of the individual Q due to its decision-making on joint actions with larger action spaces and the limited data collected in each training iteration compared to the parameter-shared individual parts. To improve the overall learning efficiency, we adopt the approach of adjusting the learning rates and the dynamic learning rate [42] to balance the training of individual Q networks for decision making with the acquisition of global knowledge, while accelerating the convergence of the global Q network.
Figure 5: The training process of the comparative experiment. It can be seen from the figure that the ddpg-type method performs the worst. QCOMBO achieves the highest reward effect at the beginning of training, which proves the effectiveness of global decision-making. On this basis, our MQLC method effectively improves the performance due to additional observation acquisition and feature extraction, and reaches the highest level compared with other models in the later stage of training.
@@
IV-B Baseline methods and Evaluation Metrics
Our experiments are conducted in B-GAP[39] environment. We select the following baseline models for comparative experiments in order to fully validate the effectiveness of our proposed method:
  • Random adopt random actions as the agent’s decision.
  • DDPG[29] combines the benefits of DQN (Deep Q-Network) and Policy Gradient methods, handling continuous action spaces effectively.
  • DDPG-MBE[43] enhances DDPG by generating additional training data through simulated rollouts.
  • DQN[44] is an algorithm that combines Q-learning with neural networks. It uses a neural network to approximate the Q-value function, which estimates the expected rewards for state-action pairs.
  • Double DQN[45, 46] uses two networks to decouple action selection from action evaluation, leading to more accurate Q-value estimates which results in improved stability and performance in training.
  • D3QN[47, 48] combines Dueling Network Architectures with Double DQN. It separates the estimation of the state value and the advantage for each action, which helps in better learning of state values and more accurate Q-value estimates. In the experiment, we set its input as state to ensure the fairness of the comparison experiment.
  • QCOMBO[30] is a multi-agent reinforcement learning algorithm for optimizing traffic signal networks. It combines independent and centralized learning using a consistency regularizer to align individual and global action-value functions. We adapt it to the lane-changing decision environment.
Among the above models, the architecture and design of DQN and its variants Double DQN and D3QN are all from the original B-GAP[39] environment. Their parameters also adopt the original parameters set by the environment.
@@ IV-B Baseline methods and Evaluation Metrics
As for the evaluation metrics, we emphasise the consideration of the reward component in the evaluation metrics, as rewards serve as the only feedback signal during reinforcement learning training. They directly reflect the quality of the actions taken by the intelligent agent. The efficiency and safety of lane-changing decisions are also the key objectives we consider. The average speed of agents provides a clear understanding of the benefits of their decisions and can also indirectly reflect the overall traffic efficiency. We therefore choose it as a measure of efficiency. And in the simulation environment, collisions lead to the termination of episodes before the maximum episode length is reached. The longer the episode length is, the safer the lane-changing decision made by the agent is. As a result, the average episode length serves as an indicator for evaluating the safety of the models.
IV-C Result of comparative experiments
@@
To demonstrate the superiority of the proposed MQLC in decision making, we conduct comparative experiments. Comparative experiments were carried out in the three traffic environments shown in Table IV. The “Sparse” mode in the table represents a low-density traffic pattern where there are two agents and 8 surrounding vehicles; the “Normal” mode represents a medium-density traffic pattern with 3 agents and 15 surrounding vehicles; and the “Dense” mode represents a high-density traffic pattern with 5 agents and 30 surrounding vehicles.
Figure 5 shows the training process of the comparative experiment. From the figure, we can clearly see that the DDPG and DDPG-MBE methods show the worst performance, which proves that this type of method is not suitable for discrete action spaces. Compared with the benchmark DQN, Double DQN and D3QN models, QCOMBO shows outstanding performance. This is because the global decision-making guidance feature of QCOMBO allows the agent to have more considerations for the environment, thereby improving the overall lane-changing effect. On this basis, the MQLC model proposed in this paper optimizes observation acquisition and feature extraction, innovatively adopts a global decision-making method, and achieves the best effect in the later stage of training. This figure further supplements the effectiveness of the method proposed in this paper.
Table V shows the results of the comparative experiments in terms of safety, efficiency and reward metrics. From Table V, we can draw the following conclusions:
(1) Comparing the “MQLC” row in Table V with the “DDPG” and “DDPG-MBE” rows, it can be seen that both models lag far behind our method in terms of average episode duration and total reward indicators. For example, MQLC outperforms DDPG-MBE by $305.9\%$ in terms of “Reward” in the “Normal” traffic scenario. The difference in this metric between the two methods is obvious. The reason why the DDPG method lags behind is that it is proposed to solve the problem of continuous action space, while the lane change decision task belongs to the discrete action space problem with fewer discrete points. The large difference in the action space between the two leads to inadaptability. This indicates that even the most advanced multi-agent collaboration method may not be applicable to lane-changing decision tasks due to the particularity of the decision-making environment and settings.
@@
(2) Comparing the “MQLC” row in Table V with the “DQN”, “Double DQN” and “D3QN” rows, it can be observed that MQLC achieves the best performance among the reinforcement learning methods. In the “Normal” and “Dense” traffic scenarios, MQLC achieves the highest average episode length, the highest speed, and the optimal reward. For instance, MQLC outperforms D3QN by $5.6\%$ and $30.5\%$ in terms of “Reward” in the “Normal” and “Dense” traffic scenarios. However, in the “Sparse” scenario, all models show similar performance and MQLC’s performance is not evidently better than the other models. This is because MQLC’s primary innovation lies in the coordination between multiple intelligent agents. When there are relatively few agents, MQLC does not show a clear advantage. However, as the number of agents increases, MQLC can optimise the overall traffic flow by using additional feature extraction and coordinated decision making, balancing individual and global interests, and thus achieving optimisation of the overall traffic flow.
@@
(3) Comparing the “MQLC” row with the “QCOMBO” row shows that MQLC outperforms QCOMBO in various scenarios in general. In terms of “Length” which evaluates the safety, MQLC outperforms QCOMBO by $35.2\%$, $19.6\%$ and $11.8\%$ in the “Sparse”, “Normal” and “Dense” traffic scenarios. This is because QCOMBO only considers bringing individual network estimates closer to the global network during training, but continues to use individual decisions during actual decision making, and is unable to overcome the limitations of information acquisition and decision coordination. In contrast, MQLC fully exploits the advantages of CVIS, uses state information to understand the overall traffic situation, and effectively coordinates during decision making by using individual agents to support global decision making. As a result, MQLC achieves significantly better lane-changing decision performance than QCOMBO.
@@
(4) Comparing the three performance metrics across various scenarios, MQLC consistently outperforms other models in terms of “Length” and “Reward” in nearly all cases. However, except for the most congested “Dense” scenario, there isn’t a significant advantage in the “Speed” metric in the other two scenarios. These results suggest that MQLC places a higher emphasis on prioritizing the safety of all agents’ actions through collaborative efforts, thereby avoiding collision penalties and ultimately achieving superior lane-changing decisions, rather than solely focusing on efficiency enhancement.
TABLE VI: The effect of architecture of intent prediction network
Architecture Episodes length Average speed Total reward
LSTM 32.87 24.30 69.49
GCN+GRU 35.08 24.34 71.22
IV-D Design of the intent prediction network
In section III-A, we introduced the design of the intention prediction part in detail, that is, to expand the observation by predicting the trajectories of vehicles around the agent to improve the ability of the agent to collaborate. In this section, we analyze the specific design of the trajectory prediction network in table I. For comparison, we implemented the LSTM-based intention prediction network under “normal” traffic density conditions while keeping other conditions the same. The experimental results are shown in table VI.
From the analysis, we can see that the final lane-changing effect of the intention prediction network based on GCN combined with GRU adopted in this paper is much better than that of the LSTM-based model in terms of episodes’ average length and comprehensive reward. This is because in the lane-changing scenario, the interaction between vehicles can be perfectly represented by the topological structure. The GCN-based model can accurately capture the interaction relationship between vehicles in the lane-changing decision, and then use the more computationally efficient GRU model for time series prediction, more accurately predicting the future trajectory of surrounding vehicles and providing a reference for decision-making. This result also proves the superiority of the architecture selected in this paper.
TABLE VII: The effect of $\lambda$ in performance